OpenStudy (anonymous):

list all subsets of {1,2}

OpenStudy (akashdeepdeb):

WELCOME TO OPENSTUDY!! :D

OpenStudy (akashdeepdeb):

All subsets have the set and null set in them. Subsets of {1,2} then would be.. {1},{2},{1,2},() Understood this? :)

OpenStudy (anonymous):

Should subset {2,1} be included? Book shows it is....why?

OpenStudy (goformit100):

"Welcome to OpenStudy. I can answer your questions or guide you. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (asnaseer):

if you have two subsets A and B, the B would be a subset of A if every member of B is also a member of A. agreed?

OpenStudy (anonymous):

0 is an element \[But null set expressed as \left\{ \right\} or \phi \] is a subset of every set.

OpenStudy (anonymous):

\[\left\{ 1,2 \right\} and \left\{ 2,1 \right\} are same.\]